# Quasinormal modes of Schwarzschild-de Sitter black holes in semi-open systems

### v1

@@ II Sets up for Schwarzschild-de Sitter black holes
where the function $f(r)$ has been parameterized by the event horizon $r_{e}$ and the cosmological horizon $r_{c}$, and $r_{n}$ is the third real root of $f(r)$ with $r_{n}=-(r_{e}+r_{c})<0$. From Eq. (1), one can directly derive the relation between the cosmological constant $\Lambda>0$ and the black hole mass $M$ and horizons via Konoplya:2022xid; Zhou:2025xta
@@
Here, we consider three kinds of potentials whose expressions are Zhidenko:2003wq; Arnaudo:2025kit
@@
where $s=0,1,2$ indicate scalar, electromagnetic and gravitational axial perturbations, respectively, and $\delta_{s,2}$, $\delta_{s,0}$ stand for the Kronecker delta. Regarding the scalar perturbation $s=0$, it should be noted that the case discussed here is the conformally coupled massless scalar perturbation Crispino:2013pya, $(\square-R/6)\Phi=0$, where for the Schwarzschild-de Sitter black hole the scalar curvature is satisfied with $R=4\Lambda$. In the case of other types of perturbations, such as the minimally coupled massless scalar, the corresponding differential equation does not simplify to a Heun’s equation. Mathematically, for the minimally coupled massless scalar, although the infinity point $r=\infty$ is a regular singular point, the difference between the two associated indicial exponents is $3$. One can use the so-called Liouville invariant to demonstrate that, with this specific structure, this regular singular point can not be removed.
Like the Pöschl-Teller potential Solidoro:2024yxi, the analytical solution for the potential (5) of the Schwarzschild-de Sitter black hole can be derived. From Appendix A, at the event horizon $z=0$, one finds the general local solution $\Psi(r)$ which can be written as
@@
The symbol $Hl$ corresponds to the Heun function, which is a built-in function written “HeunG” in Mathematica. At the cosmological horizon $z=1$, one finds the general local solution $\Psi(r)$ can be written as
@@
$D_{11}=\frac{W[u_{01},u_{12}]}{W[u_{11},u_{12}]}\,,\quad D_{12}=\frac{W[u_{01},u_{11}]}{W[u_{12},u_{11}]}\,,\quad D_{21}=\frac{W[u_{02},u_{12}]}{W[u_{11},u_{12}]}\,,\quad D_{22}=\frac{W[u_{02},u_{11}]}{W[u_{12},u_{11}]}\,$, (15)
The four functions $u_{01}(z)$, $u_{02}(z)$, $u_{11}(z)$ and $u_{12}(z)$ are from Eqs. (A) and Eqs. (A). It is not difficult to prove that the ratio of the Wronskian is independent of $z$ from which the connection coefficients in Eqs. (13) and Eqs. (15) are independent of $z$. Therefore, in the evalution of the above Wronskians, one can choose a value of $z$ located in both convergence circles at $z=0$ and $z=1$. In practice, the middle ponit $z_{0}=1/2$ is good point to achieve this purpose.
@@
The detailed expressions of $A_{\text{in}}(\omega)$, $A_{\text{out}}(\omega)$, $B_{\text{in}}(\omega)$ and $B_{\text{out}}(\omega)$ are found in Appendix B. Consider the semi-open problem in which a wall of reflectivity $\mathcal{K}(\omega)$ is located at $x_{0}\ll 0$ such that $V_{s}(x_{0})$ is approximated as zero, and the solution $\Psi_{s}(x)$ at $x=x_{0}$ is a linear combination of ingoing and outgoing waves $\mathrm{e}^{\pm\mathrm{i}\omega x}$. In other words, the solution of Eq. (3) obey the following boundary condition Berti:2025hly; Macedo:2018yoi; Solidoro:2024yxi; Rosato:2025byu; Herdeiro:2019fps
@@
From the property of being constant for the Wronskians of two solutions of Eq. (3), one can write the following three useful relations between the coefficients $A_{\text{in}}$, $A_{\text{out}}$, $B_{\text{in}}$, $B_{\text{out}}$, $S_{\text{in}}$, and $S_{\text{out}}$:
@@ III Quasinormal modes
In this section, we will study the QNM spectra for the existence of a wall of reflectivity $\mathcal{K}(\omega)$ in the Schwarzschild-de Sitter black hole. Here, we always focus on the axial gravitational perturbation, i.e., $s=2$, and the angular momentum parameter $\ell$ is set to be $2$. Perturbations of other spins satisfy the same master equation with similar effective potentials, consequently, the qualitative conclusions drawn here should remain valid for them. For the sake of simplicity, although $\mathcal{K}(\omega)$ may depend on $\omega$, in this section, we focus on the reflectivity that does not depend on the frequency. At this time, one can write $\mathcal{K}(\omega)$ as $\mathcal{K}\in\mathbb{C}$ with $|\mathcal{K}|\leq 1$. This condition allow us to vary the amount of the absorption for the wall. To be more specific, for $\mathcal{K}=1$, the wall is fully reflective, namely the Neumann boundary condition, while for $\mathcal{K}=0$, it recovers open boundary condition. Note that for $\mathcal{K}=-1$, we will get the Dirichlet boundary condition. More importantly, as $\mathcal{K}$ has an imaginary part (Note that this case will be studied associated with the EPs in Sec. V), such wall is going to lead the Robin boundary condition Solidoro:2024yxi; Herdeiro:2019fps.
@@
We are concerned with the most common scenario where $\mathcal{K}\in\mathbb{R}$ satisfies $0\leq\mathcal{K}\leq 1$. The QNM spectra of the semi-open system for the SdS black hole are obtained by numerically solving $S_{\text{in}}(\omega)=0$, namely QNM spectra are the zeros of the function $S_{\text{in}}(\omega)$. For a non-zero reflectivity, one should expect the QNM spectra will have a deviation from those of the original open problem. The problem of finding the zeros of nonlinear functions usually requires reliable iterative initial values, and Ref. Wu:2025sbq provides a practical and feasible way to find them. However, in order to capture the essence of the impact of reflectivity on QNM spectra, we consider two scenarios, namely the near extreme SdS case and the near Schwarzschild case. For these two cases, one can use the QNM spectra results from near extreme SdS case and the Schwarzschild case as the iterative initial values for getting the zeros of the function $S_{\text{in}}(\omega)$ within $\mathcal{K}=0$. For the near extremal SdS black hole, which is defined as the spacetime for which the cosmological horizon $r_{c}$ is very close to the black hole horizon $r_{e}$, i.e. $r_{e}/r_{c}\approx 1$, the analytical expressions of the QNM spectra have been found in Cardoso:2003sw; Konoplya:2004uk; Jansen:2017oag. For $s=2$, the expression is given by
@@
$\omega_{n}=\kappa_{e}\Bigg[-\Big(n+\frac{1}{2}\Big)+\sqrt{(\ell+2)(\ell-1)-\frac{1}{4}}\Bigg]\,,\quad n=0,1,\cdots\,.$ (26)
@@
The simple form of the QNM spectra in extreme SdS black holes is attributed to the behavior of the effective potential approximating the Pöshl-Teller potential Ferrari:1984zz. For the Schwarzschild black hole, one can find QNM spectra in Berti:2009kk; Mamani:2022akq; Leaver:1985ax.
@@
In Fig. 1 (near extreme case), we display their continuous migrations in the complex plane as $\mathcal{K}$ is varied, fixing $r_{e}=1$ and $r_{c}=1.1$, where in Fig. 1, the reflectivity wall is located at $x_{0}=-100$ while in Fig. 1 the reflectivity wall is located at $x_{0}=-150$. A total of $10$ mode migrations are presented. The mark $\star$ represents the QNM spectra of the open system, i.e., $\mathcal{K}=0$. As $\mathcal{K}$ is increased, the modes migrate down toward the real line. Intuitively, it can be identified that three qualitatively different types of QNMs. Some modes migrate to the real axis (in fact, the imaginary parts of them are very small), for example $n=0$, $n=3$, $n=5$ in Fig. 1 with $n=1$, $n=3$, $n=5$, $n=8$ in Fig. 1. These modes are called the quasibound states (QBS), and they have long lifetimes. Some modes, although also gradually approaching the real axis, do not have an imaginary part that is entirely zero when $\mathcal{K}=1$ for example $n=1$, $n=2$, $n=4$, $n=6$, $n=8$ in Fig. 1 with $n=0$, $n=2$, $n=4$, $n=6$, $n=7$, $n=9$ in Fig. 1. The third type of mode will migrate to the imaginary axis as $\mathcal{K}$ increases, becoming $\text{Re}(\omega)=0$, for example $n=7$, $n=9$ in Fig. 1. Although the aforementioned phenomenon has been explained in Solidoro:2024yxi, we have observed a similar phenomenon in the Schwarzschild-de Sitter black hole. Furthermore, comparing the two panels in Fig. 1, we can also conclude that the closer $x_{0}$ is to the event horizon, the closer the modes $(n\geq 10)$ are to the real axis when $\mathcal{K}=1$. It can be also concluded that for Fig. 1, those modes that collide with the imaginary axis occur at higher order overtones, although we have not shown. Additionally, as $x_{0}$ decreases, the instability of the mode spectrum increases, meaning that a smaller $\mathcal{K}$ can cause a significant shift in the spectrum.
@@
Figure 2: The migrations of the QNM spectra with $\mathcal{K}\in[0,1]$ varying in the complex plane for the first 77 modes, where the parameters of left panel are chosen $r_{e}=0$, $r_{c}=100$, and $x_{0}=-20$ while the parameters of right panel are chosen $r_{e}=0$, $r_{c}=100$, and $x_{0}=-30$. Three subfigures above are the zoom-in near the imaginary axis.
@@ IV Greybody factors
Greybody factors encode the frequency-dependent transmission of perturbations across the effective potential barrier of a black hole spacetime. They enter directly in the computation of Hawking fluxes and energy spectra. Unlike individual QNM spectra, they often behave as remarkably robust observables Oshita:2023cjz; Rosato:2024arw; Oshita:2024fzf; Ianniccari:2024ysv; Wu:2024ldo; Kyutoku:2022gbr; Konoplya:2025ixm; Wu:2025sbq; Xie:2025jbr. The robustness of the greybody factors indicates that ringdown waveform built directly from the greybody factors may offer a better modelling strategy in future gravitational wave studies. Motivated by possible quantum gravity effects, a variety of proposals for horizon scale structure (ranging from firewalls Bousso:2025udh and fuzzballs Mathur:2008kg to more general exotic compact objects Mark:2017dnq) suggest that the classical picture of a perfectly absorbing event horizon may break down at the would-be horizon. We therefore consider a semi-open system, in which the near horizon region is modelled by a partially reflecting surface.
In this section, we focus on the greybody factors (transmission coefficients) of the Schwarzschild-de Sitter black hole with a semi-open condition to account for some possible quantum corrections near the event horizon. We consider a setup in which a partially reflecting surface is introduced just outside the black hole horizon. The wall is placed at a fixed tortoise coordinate $x_{0}\ll 0$, in a region where the effective potential can be approximated as $V_{s}(x_{0})\simeq 0$. Two representative reflection models are considered: a frequency-independent (constant) reflectivity Mark:2017dnq and a Boltzmann-type reflectivity Wang:2019rcf; Oshita:2019sat. Under the boundary conditions Eq. (24) and the definition of greybody factor Rosato:2025byu,
@@
Figure 3: Greybody factor $\Gamma(\omega)$ for axial gravitational $(s=2)$ perturbations of the Schwarzschild-de Sitter black hole with a frequency-independent reflectivity $\mathcal{K}(\omega)$. In both panels the reflective wall is placed at $x_{0}=-30$, and the curves with different colors correspond to $\mathcal{K}(\omega)=0,\,0.25,\,0.50,\,0.75,$ and $1$. The left panel shows the near-extremal case with $r_{e}=1$ and $r_{c}=1.1$, and the right panel shows the nearly Schwarzschild case with $r_{e}=1$ and $r_{c}=100$.
@@
The origin of these resonances is that, when a reflective surface is introduced near the horizon, the echoes reflected back from the wall interfere with the incident wave and thus give rise to the resonant behaviour Solidoro:2024yxi. The number and positions of the resonant peaks are entirely determined by the effective distance $L_{\rm eff}$ between the wall position and the effective potential Solidoro:2024yxi; Rosato:2025byu; Rosato:2025lxb. When the wall position $x_{0}$ is fixed and $r_{c}$ is increased, or when $r_{c}$ is fixed and the reflective wall $x_{0}$ is moved closer to the horizon, $L_{\rm eff}$ increases in both cases, and the number of resonant peaks correspondingly increases. The right panel of Fig. 3 clearly shows that, when $L_{\rm eff}$ is fixed, the number and positions of the resonant peaks are almost the same for different values of $\mathcal{K}$, and only the peak amplitudes change, being correlated with $\mathcal{K}$.

### v2

@@ II Sets up for Schwarzschild-de Sitter black holes
where the function $f(r)$ has been parameterized by the event horizon $r_{e}$ and the cosmological horizon $r_{c}$, and $r_{n}$ is the third real root of $f(r)$ with $r_{n}=-(r_{e}+r_{c})<0$. From Eq. (1), one can directly derive the relation between the cosmological constant $\Lambda>0$ and the black hole mass $M$ and horizons via Konoplya and Zhidenko (2022); Zhou and Panosso Macedo (2025)
@@
Here, we consider three kinds of potentials whose expressions are Zhidenko (2004); Arnaudo and Withers (2025)
@@
where $s=0,1,2$ indicate scalar, electromagnetic and gravitational axial perturbations, respectively, and $\delta_{s,2}$, $\delta_{s,0}$ stand for the Kronecker delta. Regarding the scalar perturbation $s=0$, it should be noted that the case discussed here is the conformally coupled massless scalar perturbation Crispino et al. (2013), $(\square-R/6)\Phi=0$, where for the Schwarzschild-de Sitter black hole the scalar curvature is satisfied with $R=4\Lambda$. In the case of other types of perturbations, such as the minimally coupled massless scalar, the corresponding differential equation does not simplify to a Heun’s equation. Mathematically, for the minimally coupled massless scalar, although the infinity point $r=\infty$ is a regular singular point, the difference between the two associated indicial exponents is $3$. One can use the so-called Liouville invariant to demonstrate that, with this specific structure, this regular singular point can not be removed.
Like the Pöschl-Teller potential Solidoro et al. (2025), the analytical solution for the potential (5) of the Schwarzschild-de Sitter black hole can be derived. From Appendix A, at the event horizon $z=0$, one finds the general local solution $\Psi(r)$ which can be written as
@@
The symbol $Hl$ standing for “Heun-local” corresponds to the Heun function, which is a built-in function written “HeunG” in Mathematica. At the cosmological horizon $z=1$, one finds the general local solution $\Psi(r)$ can be written as
@@
$D_{11}=\frac{W[u_{01},u_{12}]}{W[u_{11},u_{12}]}\,,\quad D_{12}=\frac{W[u_{01},u_{11}]}{W[u_{12},u_{11}]}\,,\quad D_{21}=\frac{W[u_{02},u_{12}]}{W[u_{11},u_{12}]}\,,\quad D_{22}=\frac{W[u_{02},u_{11}]}{W[u_{12},u_{11}]}\,.$ (15)
The four functions $u_{01}(z)$, $u_{02}(z)$, $u_{11}(z)$ and $u_{12}(z)$ are from Eqs. (A) and Eqs. (A) in Appendix A. It is not difficult to prove that the ratio of the Wronskian is independent of $z$ from which the connection coefficients in Eqs. (13) and Eqs. (15) are independent of $z$. Therefore, in the evalution of the above Wronskians, one can choose a value of $z$ located in both convergence circles at $z=0$ and $z=1$. In practice, the middle ponit $z_{0}=1/2$ is good point to achieve this purpose.
@@
The detailed expressions of $A_{\text{in}}(\omega)$, $A_{\text{out}}(\omega)$, $B_{\text{in}}(\omega)$ and $B_{\text{out}}(\omega)$ are found in Appendix B. Consider the semi-open problem in which a wall of reflectivity $\mathcal{K}(\omega)$ is located at $x_{0}\ll 0$ such that $V_{s}(x_{0})$ is approximated as zero, and the solution $\Psi_{s}(x)$ at $x=x_{0}$ is a linear combination of ingoing and outgoing waves $\mathrm{e}^{\pm\mathrm{i}\omega x}$. To give a clear description of the proximity between the wall and the event horizon, we convert the wall position $x_{0}$ into the corresponding areal radial coordinate $r_{0}$ and use $\Delta\equiv(r_{0}-r_{e})/r_{e}$ to portray. For such a situation, the solution of Eq. (3) obeys the following boundary condition Berti and others (2025); Macedo et al. (2018); Solidoro et al. (2025); Rosato et al. (2025b); Herdeiro and Santos (2019)
@@
From the property of being constant for the Wronskians of two solutions of Eq. (3), one can write the following three useful relations between the coefficients $A_{\text{in}}(\omega)$, $A_{\text{out}}(\omega)$, $B_{\text{in}}(\omega)$, $B_{\text{out}}(\omega)$, $S_{\text{in}}(\omega)$, and $S_{\text{out}}(\omega)$:
@@ III Quasinormal modes
In this section, we will study the QNM spectra for the existence of a wall of reflectivity $\mathcal{K}(\omega)$ in the Schwarzschild-de Sitter black hole. Here, we always focus on the axial gravitational perturbation, i.e., $s=2$, and the angular momentum parameter $\ell$ is set to be $2$. Perturbations of other spins satisfy the same master equation with similar effective potentials, consequently, the qualitative conclusions drawn here should remain valid for them (One can find a special case of QNMs for $s=0$ in Appendix C.). For the sake of simplicity, although $\mathcal{K}(\omega)$ may depend on $\omega$, in this section, we focus on the reflectivity that does not depend on the frequency. At this time, one can write $\mathcal{K}(\omega)$ as $\mathcal{K}\in\mathbb{C}$ with $|\mathcal{K}|\leq 1$. This condition allow us to vary the amount of the absorption for the wall. To be more specific, for $\mathcal{K}=1$, the wall is fully reflective, namely the Neumann boundary condition, while for $\mathcal{K}=0$, it recovers open boundary condition. Note that for $\mathcal{K}=-1$, we will get the Dirichlet boundary condition. More importantly, as $\mathcal{K}$ has an imaginary part (Note that this case will be studied associated with the EPs in Sec. V), such wall is going to lead the Robin boundary condition Solidoro et al. (2025); Herdeiro and Santos (2019).
@@
We are concerned with the most common scenario where $\mathcal{K}\in\mathbb{R}$ satisfies $0\leq\mathcal{K}\leq 1$. The QNM spectra of the semi-open system for the SdS black hole are obtained by numerically solving $S_{\text{in}}(\omega)=0$, namely QNM spectra are the zeros of the function $S_{\text{in}}(\omega)$. For a non-zero reflectivity, one should expect the QNM spectra will have a deviation from those of the original open problem. The problem of finding the zeros of nonlinear functions usually requires reliable iterative initial values, and Ref. Wu et al. (2025b) provides a practical and feasible way to find them. However, in order to capture the essence of the impact of reflectivity on QNM spectra, we consider two scenarios, namely the near extreme SdS case and the near Schwarzschild case. For these two cases, one can use the QNM spectra results from near extreme SdS case and the Schwarzschild case as the iterative initial values for getting the zeros of the function $S_{\text{in}}(\omega)$ within $\mathcal{K}=0$. For the near extremal SdS black hole, which is defined as the spacetime for which the cosmological horizon $r_{c}$ is very close to the black hole horizon $r_{e}$, i.e., $r_{e}/r_{c}\approx 1$, the analytical expressions of the QNM spectra have been found in Cardoso and Lemos (2003); Konoplya and Zhidenko (2004); Jansen (2017). For $s=2$, the expression is given by
@@
$\omega_{n}=\kappa_{e}\Bigg[-\Big(n+\frac{1}{2}\Big)\mathrm{i}+\sqrt{(\ell+2)(\ell-1)-\frac{1}{4}}\Bigg]\,,\quad n=0,1,\cdots\,$, (26)
@@
where $\kappa_{e}$ is the surface gravity at the event horizon. The simple form of the QNM spectra in extreme SdS black holes is attributed to the behavior of the effective potential approximating the Pöshl-Teller potential Ferrari and Mashhoon (1984). For the Schwarzschild black hole, one can find QNM spectra in Berti et al. (2009); Mamani et al. (2022); Leaver (1985).
@@
In Fig. 1 (near extreme case), we display their continuous migrations in the complex plane as $\mathcal{K}$ is varied, fixing $r_{e}=1$ and $r_{c}=1.1$, where in Fig. 1, the reflectivity wall is located at $x_{0}=-100$ while in Fig. 1 the reflectivity wall is located at $x_{0}=-150$. A total of $10$ mode migrations are presented. The mark $\star$ represents the QNM spectra of the open system, i.e., $\mathcal{K}=0$. As $\mathcal{K}$ is increased, the modes migrate down toward the real line. Intuitively, it can be identified that three qualitatively different types of QNMs, which is also emphasized and explained in recent work Solidoro et al. (2025). First type of modes migrate to the real axis (in fact, the imaginary parts of them are very small), for example $n=0$, $n=3$, $n=5$ in Fig. 1 with $n=1$, $n=3$, $n=5$, $n=8$ in Fig. 1. These modes become the quasibound states (QBS) which have long lifetimes. This behavior is naturally understood as follow. The real part $\text{Re}(\omega_{0})$ serves as a practical threshold for traversing the potential barrier, where $\omega_{0}$ is the original fundamental QNM. These modes with $0<\text{Re}(\omega)<\text{Re}(\omega_{0})$ lie effectively below this threshold, cannot propagate across the photon-sphere region, and are therefore trapped within the wall–barrier cavity, escaping only via tunneling through the effective potential barrier, which naturally yields very small imaginary parts and long lifetimes. The second branch of modes, although also gradually approaching the real axis, do not have an imaginary part that is entirely zero when $\mathcal{K}=1$ for example $n=1$, $n=2$, $n=4$, $n=6$, $n=8$ in Fig. 1 with $n=0$, $n=2$, $n=4$, $n=6$, $n=7$, $n=9$ in Fig. 1. Which correspond to the over-barrier regime, $\text{Re}(\omega)>\text{Re}(\omega_{0})$: they exceed the effective energy threshold for traversing the photon-sphere potential barrier, therefore they effectively pass over the barrier, experience only weak reflection of potential and escape to infinity. Consequently, they are more strongly damped, retaining a finite decay rate and can be identified with the “new overtones” of the semi-open system. Last, the third type of mode will migrate to the imaginary axis as $\mathcal{K}$ increases, becoming $\text{Re}(\omega)=0$, for example $n=7$, $n=9$ in Fig. 1. These modes are caused by the presence of attractors from the perspective of dynamical system Torres and Dolan (2026). Although they study a localized modification of the effective potential of the form $\epsilon\delta(x-x_{0})$, rather than the changing the boundary condition considered here, these two approaches are physically equivalent in the relevant limit: as $\epsilon$ increases, the $\epsilon\delta(x-x_{0})$ modification acts increasingly like an effective reflective wall. They identify attractors on the imaginary axis, which naturally accounts for the emergence of purely damped modes. Furthermore, comparing the two panels in Fig. 1, we can also conclude that the closer $x_{0}$ is to the event horizon, the closer the modes $(n\geq 10)$ are to the real axis when $\mathcal{K}=1$. It can be also concluded that for Fig. 1, those modes that collide with the imaginary axis occur at higher order overtones, although we have not shown. Additionally, as $x_{0}$ decreases, the instability of the mode spectrum increases, meaning that a smaller $\mathcal{K}$ can cause a significant shift in the spectrum.
@@
Figure 2: The migrations of the QNM spectra with $\mathcal{K}\in[0,1]$ varying in the complex plane for the first 77 modes, where the parameters of panel (a) are chosen $r_{e}=1$, $r_{c}=100$, and $x_{0}=-20$ $(\Delta=4.6490692\times 10^{-8})$ while the parameters of panel (b) are chosen $r_{e}=1$, $r_{c}=100$, and $x_{0}=-30$ $(\Delta=2.1169523\times 10^{-12})$. Three subfigures above are the zoom-in near the imaginary axis.
@@ IV Greybody factors
Greybody factors encode the frequency-dependent transmission of perturbations across the effective potential barrier of a black hole spacetime. They enter directly in the computation of Hawking fluxes and energy spectra. Unlike individual QNM spectra, they often behave as remarkably robust observables Oshita (2024); Rosato et al. (2024); Oshita et al. (2024); Ianniccari et al. (2024); Wu et al. (2025a); Kyutoku et al. (2023); Konoplya and Pappas (2025); Wu et al. (2025b); Xie et al. (2025). The robustness of the greybody factors indicates that ringdown waveform built directly from the greybody factors may offer a better modelling strategy in future gravitational wave studies. Motivated by possible quantum gravity effects, a variety of proposals for horizon scale structure (ranging from firewalls Bousso (2025) and fuzzballs Mathur (2010) to more general exotic compact objects Mark et al. (2017)) suggest that the classical picture of a perfectly absorbing event horizon may break down at the would-be horizon. We therefore consider a semi-open system, in which the near horizon region is modelled by a partially reflecting surface.
In this section, we focus on the greybody factors (transmission coefficients) of the Schwarzschild-de Sitter black hole with a semi-open condition to account for some possible quantum corrections near the event horizon. We consider a setup in which a partially reflecting surface is introduced just outside the black hole horizon. The wall is placed at a fixed tortoise coordinate $x_{0}\ll 0$, in a region where the effective potential can be approximated as $V_{s}(x_{0})\simeq 0$. Two representative reflection models are considered: a frequency-independent (constant) reflectivity Mark et al. (2017) and a Boltzmann-type reflectivity Wang et al. (2020); Oshita et al. (2020). Under the boundary condition (24) and the definition of greybody factor Rosato et al. (2025b),
@@
Figure 3: Greybody factor $\Gamma(\omega)$ for axial gravitational $(s=2)$ perturbations of the Schwarzschild-de Sitter black hole with a frequency-independent reflectivity $\mathcal{K}(\omega)$. In both panels the reflective wall is placed at $x_{0}=-30$ ($\Delta=4.5730021\times 10^{-3}$ and $\Delta=2.1169523\times 10^{-12}$), and the curves with different colors correspond to $\mathcal{K}(\omega)=0,\,0.25,\,0.50,\,0.75,$ and $1$. Panel (a) shows the near-extremal case with $r_{e}=1$ and $r_{c}=1.1$, and panel (b) shows the nearly Schwarzschild case with $r_{e}=1$ and $r_{c}=100$.
@@
The origin of these resonances is that, when a reflective surface is introduced near the horizon, the echoes reflected back from the wall interfere with the incident wave and thus give rise to the resonant behaviour Solidoro et al. (2025). The number and positions of the resonant peaks are entirely determined by the effective distance $L_{\rm eff}$ between the wall position and the effective potential Solidoro et al. (2025); Rosato et al. (2025b, a). When the wall position $x_{0}$ is fixed and $r_{c}$ is increased, or when $r_{c}$ is fixed and the reflective wall $x_{0}$ is moved closer to the horizon, $L_{\rm eff}$ increases in both cases, and the number of resonant peaks correspondingly increases. The right panel of Fig. 3 clearly shows that, when $L_{\rm eff}$ is fixed, the number and positions of the resonant peaks are almost the same for different values of $\mathcal{K}$, and only the peak amplitudes change, being correlated with $\mathcal{K}$.